OpenStudy (anonymous):

Find the slope of the line described by: x - 6*y = -8

OpenStudy (anonymous):

x- 6y = -8 -x -x -6y = -1x - 8 divide by -6 y = 1/6x + 4/3

OpenStudy (anonymous):

slope will be rise 1 run 6

OpenStudy (anonymous):

so 1/6

OpenStudy (anonymous):

wow thank you soo much i gave a medal